OpenStudy (anonymous):

A cylinder has its height doubled and its radius cut to one third. What is the ratio of the volumes of the modified cylinder to the original cylinder?

OpenStudy (anonymous):

do you know the formula for volume of a cylinder?

OpenStudy (anonymous):

volume = Pi r^2 h

OpenStudy (anonymous):

so if height is doubled and radius is cut to one third, new volume would be v = pi*(1/3)^2 * (2h) right?

OpenStudy (anonymous):

I'm not sure but it sounds right

OpenStudy (anonymous):

wait yeah i forgot to put something. v = pi * (r/3)^2 + 2h there that's right now. so what will be the new volume?

OpenStudy (anonymous):

I do not know.

OpenStudy (anonymous):

what is the square of 1/3?

OpenStudy (anonymous):

lol wait. i have another typo sorry. v = pi*(r/3)^2 * 2h sorry bout that

OpenStudy (anonymous):

its okay. But Idk what that is .

OpenStudy (anonymous):

(1/3)^2 = (1)^2/(3)^2

OpenStudy (anonymous):

so the both equal 1/9 ?

OpenStudy (anonymous):

what is the square of r?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

it's r^2

OpenStudy (anonymous):

square of r is r squared

OpenStudy (anonymous):

so it becomes v = pi * (1/9 r^2) * 2h what is 2 x 1/9?

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

so it becomes v = 2/9 * (pi * r^2 * h) do you notice something?

OpenStudy (anonymous):

not really. lol ..sorry I'm stupid :P

OpenStudy (anonymous):

pi*r^2*h is the original volume

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so do you know what the ratio of the modified cylinder to the original cylinder is?

OpenStudy (anonymous):

2/9 ? right

OpenStudy (anonymous):

close. the volume of modified clinder is 2/9* pi * r^2 * h. The volume of the original cylinder is pi * r^2 *h. so the ratio of modified to original would be \[\LARGE \frac{\frac29 \pi r^2 h}{\pi r^2 h}\]

OpenStudy (anonymous):

what does that give you?

OpenStudy (anonymous):

I'm not sure :S

OpenStudy (anonymous):

pi*r^2 * h cancels out \[\LARGE \frac{\frac29 \cancel{\pi r^2 h}}{\cancel{\pi r^2 h}}\] so it becomes \[\frac{\frac29}{1}\] so the ratio is 2/9 : 1

OpenStudy (anonymous):

Ohhhh. Okay. Thanks!!

OpenStudy (anonymous):

if you get confused by a problem like this use actual numbers since the answer does not depend on the actual numbers used, you will get the same answer as if you used a formula plus it is a good check

OpenStudy (anonymous):

and pick easy numbers. say the original cylinder has height 1 and radius 3 then the area is \(\pi \times 3^2=9\pi\) double the height to 2, take one third of the radius get 1 now the area is \[2\pi\] and \[\frac{2\pi}{9\pi}=\frac{2}{9}\]